1. The Creep Company has the following accounts in year-end 2020.  
  2. Kindly prepare the SFP in both report and account forms. 

 

 

Accounts Payable ₱  600,000 

Unearned Revenue 1,210,000 

Salaries Payable- current  120,000 

Owner’s Withdrawals   280,000 

Owner’s Capital 1,500,000 

Comprehensive Income  3,000,000 

 

Total Current Assets 80% of total liabilities 

 

Cash 250,000 

Inventories 480,000 

Prepaid Expenses 600,000 

Property, Plant and Equipment ? 

Accounts Receivable ? 



1
Expert's answer
2021-10-04T10:17:14-0400

Solution:

The SFP as per account and report form have been prepared as follows:

Total liabilities = 1,930,000

Total Current Assets = 80"\% \times" 1,930,000 = 1,544,000

Accounts Receivable = 1,544,000 – (250,000 + 480,000 + 600,000) = 214,000

Property, Plant and Equipment = Total Liabilities & Owners Equity – Total Current Assets = 6,710,000 – 1,544,000 = 5,166,000